Unknowns for 1H NMR Final Report 16) Chemical Formula: CaHO IR: strong peak 1720cm зн зн 2H PPM 17) Chemical Formulae: CiHs SIf someone explain how to find these unknowns..And what they are and how you got the answer. I am having a hard time really understanding NMR and IR.. Thanks in advance
Answer #1

16)   Chemical formula: C4H8O2

             Double bond equivalence or degree of unsaturation

                                 = C+1-(H/2)-(X/2)+(N/2)

        Where,

                  C: Number of Carbon atoms

                  H: Number of Hydrogen atoms

                  X: Number of Halogen atoms

                  N: Number of nitrogen atoms

Hence

DBE= 4+1-(8/2)-(0/2)-(0/2)

       = 1

Thus the compound contains either one double bond or one ring.

A strong peak at 1720 cm-1 in the IR spectrum indicate the presence of a Carbonyl group (C=O) associated with an ester moiety. C=O stretching vibrations in esters occur at 1720 – 1750 cm-1.

Now on examining 1H NMR,

  • The triplet at 1.3 ppm indicates the presence of an aliphatic –CH3 group (3H). It’s splitting pattern suggest that this –CH3 must have two similar hydrogen’s as neighbours.
  • The quartet at 2.4 ppm indicates the presence of an O=C–CH2 (2H) moiety.
  • The singlet at 3.6 ppm denotes the -O-CH3 (3H) protons

Thus the compound is

methyl propionate

17)   Chemical formula: C12H18

             Double bond equivalence or degree of unsaturation

                                 = C+1-(H/2)-(X/2)+(N/2)

        Where,

                  C: Number of Carbon atoms

                  H: Number of Hydrogen atoms

                  X: Number of Halogen atoms

                  N: Number of nitrogen atoms

Hence

DBE= 12+1-(18/2)-(0/2)-(0/2)

       = 4

Thus the compound contains an aromatic benzene ring (when DBE is 4 or above the molecule may contain an aromatic benzene ring).

Now on examining 1H NMR,

  • The triplet at 1.2 ppm indicates the presence of an aliphatic –CH3 group (3H). It’s splitting pattern suggest that this –CH3 must have two similar hydrogen’s as neighbours.
  • The multiplet at 7-7.5 ppm indicates the presence of a para disubstituted benzene ring (para substituted benzene rings have characteristic 4 four line pattern).
  • The singlet at 1.4 ppm denotes the presence of 9 aliphatic hydrogen’s associated with -C(CH3)3 moiety. The higher ppm values (usual range for -CH3 protons is 0.8 ppm) suggest that the -C(CH3)3 group is directly attatched to benzene ring.
  • A quartet at 2.7 ppm represents a –CH2– group directly attached to the benzene ring

Thus the compound is

1-(tert-butyl)-4-ethylbenzene

18)   Chemical formula: C9H12O

             Double bond equivalence or degree of unsaturation

                                 = C+1-(H/2)-(X/2)+(N/2)

        Where,

                  C: Number of Carbon atoms

                  H: Number of Hydrogen atoms

                  X: Number of Halogen atoms

                  N: Number of nitrogen atoms

Hence

DBE= 9+1-(12/2)-(0/2)-(0/2)

       = 4

Thus the compound contains an aromatic benzene ring (when DBE is 4 or above the molecule may contain an aromatic benzene ring).

A strong broad peak at 3200 cm-1 in IR spectrum indicates the presence of –OH group with intermolecular hydrogen bonding. The broadening of the band occurs due to intermolecular hydrogen bonding.

Now on examining 1H NMR,

  • The multiplet at 7-7.5 ppm indicates the presence of a mono disubstituted benzene ring (5H).
  • The triplet at 0.8 ppm indicates the presence of an aliphatic –CH3 group (3H). It’s splitting pattern suggest that this –CH3 must have two similar hydrogen’s as neighbours.
  • The multiplet (5 signals) at 1.7 ppm indicates the presence of an –CH2 (2H) moiety, surrounded by a total of 5 hydrogen’s
  • The singlet at 4.7 ppm indicates the presence of –OH group proton.
  • The triplet at 4.4 ppm denotes the –O-CH- (1H) group proton .

Thus the compound is

ОН 1-phenylpropan-1-ol